# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@ 4.1.6. Emphasizing secure practices
Key strategies to mitigate these risks include: (1) Sanitizing Generated Code: Carefully reviewing and cleaning LLM outputs to prevent vulnerabilities or unintended exposures. (2) Maintaining Data Integrity: Ensuring the confidentiality and safety of proprietary code and sensitive data. (3) Upholding Security Standards: Consistently applying established security practices to avoid compromising systems or violating regulations. By proactively addressing these issues, developers can maintain secure workflows and responsibly integrate LLM tools into their practices.
4.2. Software Engineering Tools and Designs for Developers
4.2.1. Educating developers on the benefits of LLM-powered IDE extensions
Integrating LLMs into widely-used IDEs such as Visual Studio Code (VS code) can significantly optimize development workflows and reduce friction. Extensions like GitHub Copilot, Amazon Q, and IBM watsonx are already helping developers streamline different tasks such as debugging, code generation, and documentation writing. Some participants in this study mentioned the challenges of translating LLM results into their own context, highlighting the benefits of a more streamlined user experience.
4.2.2. Tailoring answers for specific industries
Developers may benefit from LLM tools that provide tailored responses based on organization-specific or proprietary documents and data. These tools could be trained on or fine-tuned with project documentation to assist with legacy or brownfield development. Alternatively, they could use advanced AI techniques, such as retrieval-augmented generation (RAG) (Lewis et al., 2020), which would dynamically incorporate relevant artifacts like project documents, code repositories, and communication threads into the generative process to deliver more contextually-aware answers.
4.2.3. Tailoring answers for specific individuals
As a developer works over time with an LLM, contextual grounding methods can improve the models’ understanding of developers’ specific needs. Persistent context memory, for instance, could allow models to “remember” quirks and details about the user and retain them as context for future responses. For instance, an LLM could detect that a developer prioritizes attributes like performance over readability; as another example, an LLM could identify areas in which a developer frequently asks for explanations and proactively provide more detail in those areas.
4.3. Developments in LLMs and Ongoing Limitations
Since the interviews conducted during the Spring and Summer of 2023, the landscape of LLMs has notably evolved. The popular GPT-3.5 model, which was the most common model among our participants, has been replaced by GPT-4o for OpenAI’s free-tier users. For those having premium subscriptions, more advanced models, including GPT-4o, OpenAI o1, and OpenAI o1-mini, are accessible. Additionally, GitHub Copilot, the coding assistant used by two of our participants, is now also powered by GPT-4o. Table 3 provides an overview of the OpenAI and Google models’ features and performance.  [^13] Two metrics, Quality Index and Latency, are based on evaluations from Artificial Analysis (art, [n. d.]), an independent team focused on benchmarking and evaluating AI models. These metrics may not reflect the official evaluations of corresponding companies. These were selected as developers of the models discussed in this paper. Other state-of-the-art models, such as Llama by Meta and Claude by Anthropic, also demonstrate significant advancements but are outside the scope of this discussion.
Developer Model Context Window Max Output Tokens Knowledge Cut-Off Quality Index (art, [n. d.]) Latency (art, [n. d.]) Developer’s Description of Model
OpenAI GPT-4o 128k tokens 128k tokens Oct 2023 77 0.43 High-intelligence flagship model for complex, multi-step tasks. GPT-4o is cheaper and faster than GPT-4 Turbo.
GPT-4o mini 128k tokens 16k tokens Oct 2023 72 0.41 Affordable and intelligent small model for fast, lightweight tasks. GPT-4o mini is cheaper and more capable than GPT-3.5 Turbo.
o1 128k tokens 32k tokens Oct 2023 85 24.60 Reasoning model designed to solve hard problems across domains. Trained with reinforcement learning to perform complex reasoning.
o1-mini 128k tokens 64k tokens Oct 2023 82 8.93 Faster and cheaper reasoning model particularly good at coding, math, and science. Trained with reinforcement learning to perform complex reasoning.
GPT-3.5 16k tokens 4k tokens Sep 2021 53 0.38 Understands and generates natural language or code and has been optimized for chat but works well for non-chat tasks as well.
Google Gemini 1.5 Pro 2M tokens 8k tokens Sep 2024 80 1.06 Complex reasoning tasks requiring more intelligence.
Gemini 1.5 Flash 1M tokens 8k tokens Sep 2024 68 0.24 Fast and versatile performance across a diverse variety of tasks.
Table 3. Current Status of Language Models Referenced in the Paper. Quality Index represents the average performance across various evaluations of model intelligence, including benchmarks like MMLU, GPQA, and HumanEval. Latency denotes the time to the first token received after an API request, measured in seconds.
Current LLMs’ developments address a few limitations (see the blue bullet points in Figure 1) found during the interviews. For example, larger context windows now enable improved summarization and response latency has been reduced in models like GPT-4o compared to GPT-4. Issues such as hallucinations, contradictory answers, mixing up programming languages, and struggle with unstructured data have been mitigated to some extent through enhanced reasoning abilities and a larger context window.
Despite these improvements, many challenges still persist (see Challenges under different subsections in Section 3). For example, real-time browsing has been integrated into the chat interface of both GPT-4o and Gemini 1.5 to address outdated responses and lack of references. However, outputs rely heavily on static training data, making them non-referenceable. This is especially problematic in rapidly evolving information like programming libraries, where outdated information harms accuracy.
Beyond existing and future advancements, there are two considerations that are worth noting:
  1. (1)
    While flagship models like GPT-4o and o1 have addressed some limitations, smaller or locally deployable models (e.g., Llama 3.2 1B) that appeal to those prioritizing privacy or cost-efficiency continue to face challenges such as constrained context windows and reasoning capabilities.
  2. (2)
    Many limitations discussed in this paper are systemic and intrinsic to LLMs, such as their potential to impede critical learning processes if misused or their inability to replicate nuanced human interactions.
Hence, informed deployment of LLM technologies is essential to leverage their benefits while minimizing any potential harm.
5. Related Works
This paper examines the role of LLMs in software engineering, focusing on their impact on developers, the development life cycle, products developed, and societal implications. Due to the emerging nature of research in this field, we include some non-peer-reviewed studies and provide transparency through footnotes, acknowledging that, as noted by Fan et al. (Fan et al., 2023b), formal literature surveys can no longer capture all relevant work, making our review not exhaustive.
The most notable work on LLMs in the Software Engineering (SE) domain is a systematic review by Hou et al., which analyzed 395 papers on their impact, focusing on optimization techniques, applications, and potential use cases (Hou et al., 2023). In contrast, our study provides a qualitative analysis based on interviews with full-time developers.
Several interview-based studies have explored developers’ use of generative AI. Klemmer et al. found that while AI assistants like Copilot and ChatGPT were widely used for security-critical tasks, developers lacked trust in these tools and double-checked their work (Klemmer et al., 2024). Similarly, Rabani et al. reported that developers noted ChatGPT’s inaccuracy and the need for debugging (Rabani et al., 2023). Mendes et al. explored developers’ views on intelligent assistants, noting benefits like faster development and improved code but also challenges like poor accuracy and distractions (Mendes et al., 2024). Empirical studies, including those by Rasnayaka et al., assessed LLMs’ usefulness in programming projects, finding they help with code generation and debugging but highlighting a learning curve and no significant difference in software quality between AI-assisted and non-assisted teams (Rasnayaka et al., 2024).
Our study extends and corroborates these findings by exploring the broader interplay of LLMs across the four dimensions of our research questions (RQs). Accordingly, the remainder of the related work is structured around these dimensions.
5.1. People - Software Developers using LLMs
Recent studies, particularly pre-published ones, highlight the growing adoption of LLMs among professional developers. These studies, mainly empirical, include interviews and thematic analyses of developer’s written responses. Feng et al. analyzed social media posts and found that developers use ChatGPT for code debugging, interview preparation, and solving academic assignments, with fear being the predominant emotion related to code generation (Feng et al., 2023). Süße et al. identified fourteen coping patterns in a case study of AI-powered chatbots in software development, with one pattern considering the AI as a virtual colleague (Süße et al., 2023). Nam et al. found that LLMs, when used within programming environments, improve task completion by providing contextualized queries, offering a more effective alternative to simple web searches (Nam et al., 2024). Peng et al. discovered that GitHub Copilot helped developers complete tasks 55.8% faster than a control group (Peng et al., 2023), while Vaithilingam et al. noted that while Copilot did not significantly improve task completion time, it served as a useful starting point despite challenges in debugging (Vaithilingam et al., 2022). Kuhlail et al. surveyed 99 developers, finding that ChatGPT improved productivity by helping generate generic code, explain complex code, and find sources (Kuhail et al., 2024). Siddiq et al.’s study on DevGPT usage also highlighted ChatGPT’s role in helping developers understand libraries and frameworks, and engage in networking and messaging (Siddiq et al., 2024). Stack Overflow’s 2024 survey found that 63.2% of professional developers were using AI tools in the development process, with 13.5% sharing they planned to and 23.4% sharing that they did not wish to use these tools (n.a., 2024e). The Stack Overflow 2023 survey highlighted AI usage for tasks like code writing, debugging, documentation, and testing(n.a., 2023), and the 2024 survey noted that professional developers believed that AI tools could increase productivity, speed up learning, improve efficiency and code accuracy, and make workload more manageable (n.a., 2024e).
Further interview studies reveal how developers incorporate LLMs in their daily work. Pinto et al. found that LLMs, built on GPT-4, reduced repetitive tasks and helped contextualize code, though they had technical limitations like poor UI and inaccurate suggestions (Pinto et al., 2024). In a 2024 study by Kimbel et al., 17 participants across sectors reported using ChatGPT for content generation, information retrieval, brainstorming, and programming but noted challenges with accuracy (Kimbel et al., 2024) [^14] Pre-published via ResearchGate.. Coutinho et al. found that AI tools helped software professionals save and organize their time, but with reliability issues in generated content (Coutinho et al., 2024).
While these studies provide valuable insights, they fall short in addressing critical gaps. Most focus narrowly on either professional or novice developers, specific tasks, or quantitative metrics like task completion time. To date, little qualitative research has comprehensively examined the experiences of professional developers across varying experience levels, exploring not only the tasks where LLMs excel but also those where they fall short. Our study addresses this gap by investigating the nuanced ways LLMs influence different development tasks, offering a richer understanding of their practical utility and limitations.
5.2. Processes - Software Development Life Cycle Process
Emerging research into the impact of LLMs, particularly ChatGPT and GitHub Copilot, on the Software Development Life Cycle (SDLC) has identified significant effects, though more research is needed to fully understand their scope and quality. Prior to ChatGPT, earlier models like BERT were employed for specific coding tasks, such as vulnerability detection with 99.3% accuracy (Alqarni and Azim, 2022), while Text-to-Text Transfer Transformers demonstrated promise for code completion (Ciniselli et al., 2021). GitHub Copilot shifted the focus from code writing to code evaluation, with studies noting that developers often spent more time evaluating AI-generated suggestions than writing code itself (Bird et al., 2023). While Copilot provides significant value for experts, it poses risks for novices, who may struggle to identify or correct buggy or non-optimal code (Moradi Dakhel et al., 2023).
Recent studies have explored ChatGPT and GPT models across various SDLC phases. For instance, ChatGPT has proven useful in debugging (Surameery and Shakor, 2023b) and generating software requirements, though the outputs are typically less detailed than those produced by humans (Bencheikh and Höglund, 2023). It has also been employed in refining requirements (Ahmad et al., 2023). Krishna et al. demonstrated that advanced models like CodeLlama and GPT-4 could generate Software Requirement Specifications (SRS) at a level comparable to an entry-level engineer (Krishna et al., 2024) [^15] Pre-published via arXiv.. While LLMs have shown potential to assist with planning, design, implementation, and testing, they still require human supervision, particularly during the coding phase (Özpolat et al., 2023). Furthermore, Sridhara et al. found that while LLMs excel at tasks like refactoring, they struggle with more nuanced activities such as code reviews and vulnerability detection (Sridhara et al., 2023) [^16] Pre-published via arXiv..
Studies also highlight LLM contributions to software testing. Gu (Gu, 2023) and Tang (Tang et al., 2023a) found that LLMs could outperform traditional tools in test coverage, but challenges like prompt design and accuracy persist. Frameworks for evaluating LLM-generated code have been developed, including Yeo et al.’s work on prompt engineering (Yeo et al., 2024), Liu et al.’s framework for error identification (Liu et al., 2023b), and Hou et al.’s analysis of metrics like MRR, BLEU, and ROUGE to assess LLM performance in software tasks (Hou et al., 2023).
Finally, the advent of LLM-based agents marks a shift toward a more AI-enhanced SDLC. Jin et al. demonstrated that LLM-based agents, equipped with capabilities like autonomous reasoning and tool usage, handle complex tasks more efficiently than traditional LLMs (Jin et al., 2024) [^17] Pre-published via arXiv.. This shift signals a transformation in software engineering workflows, where AI roles extend from copilots to supervisors, driving new paradigms in the SDLC process (Ozkaya, 2023a, 2022).
While prior research has provided valuable insights, it remains fragmented and often focuses narrowly on specific SDLC phases or tasks. Studies rarely examine the end-to-end impact of LLMs on all SDLC activities or provide comprehensive guidance for integrating LLMs into these processes. Our research fills this gap by systematically investigating developers’ experiences using LLMs across the entire SDLC. We uniquely organize our findings according to the SDLC steps and make actionable recommendations, identifying tasks where LLMs excel and those where they under-perform. This comprehensive approach advances understanding of LLMs’ role in software development and offers practical strategies for their effective use.
5.3. Products- Artifacts
The use of LLMs in code generation has led to growing interest in understanding the quality of generated code. Studies have focused on readability (Dantas et al., 2023), complexity (Sakib et al., 2023), correctness (Liu et al., 2023b), and security (Gupta et al., 2023). For example, Nascimento et al. showed that ChatGPT outperforms beginner programmers but not experts (Nascimento et al., 2023), while Fan et al. identified shared common mistakes between human-written and generated code (Fan et al., 2023a). Stack Overflow’s 2024 Survey found that almost half of professional developers believed that AI tools were bad at handling complex tasks; additionally, developers were split on the trustworthiness of AI tools, with newer developers trusting AI accuracy more than professionals (n.a., 2024e).
Despite significant advances, much of the existing research emphasizes objective evaluations of code quality, often relying on predefined metrics or benchmarks. These studies rarely delve into developers’ subjective perceptions of the quality and trustworthiness of artifacts generated by LLMs, particularly in professional settings where these tools are actively integrated into workflows. To the best of our knowledge, limited research has explored how professional developers assess and trust the quality of artifacts produced by LLMs. Our work addresses this gap by investigating developers’ perceptions, their confidence in these outputs, and the factors influencing their trust in LLM-generated artifacts. This perspective is critical to understanding the practical integration of LLMs into professional software development and guiding improvements in LLM capabilities.
5.4. Society: LLMs in Industry and Education
5.4.1. Software Industry
LLMs are poised to significantly impact various sectors, including software development, due to their widespread adoption and diverse capabilities (Aydin and Karaarslan, 2023; Eloundou et al., 2023). A 2018 National Bureau of Economic Research report by Bessen suggests AI could reshape the labor market by replacing, shifting, or creating jobs depending on demand (Bessen, 2018). With LLMs capable of generating code, concerns have emerged about their potential to assist or replace developers. Carleton et al.’s 2021 report advocates for AI’s collaboration with software engineers to enhance productivity and reliability (Carleton et al., 2022). Kuhail et al. found that over two-thirds of developers surveyed did not foresee an immediate job security threat from AI, though many recognized a partial risk (Kuhail et al., 2024). Rashid speculated that AI could both replace some roles and open new opportunities, helping developers become more efficient (Rashid, 2023). Demirci et al. observed a 21% decrease in job postings related to coding after ChatGPT’s introduction, while jobs requiring manual labor saw less impact (Demirci et al., 2023). Additionally, research by Winter et al. showed developers prefer working alongside tools rather than having their tasks entirely replaced, suggesting that AI could support, but not fully replace, developers’ work (Winter et al., 2022).
As companies increasingly release policies on generative AI use, a 2024 interview study of 17 professionals revealed that many still lack formal guidelines for integrating AI tools like ChatGPT into their workflows (Kimbel et al., 2024).
Despite these advances, much of the existing research focuses on broad market trends, theoretical implications, or objective measures of AI’s impact. Few studies have directly interviewed developers to understand their perspectives on how LLMs are shaping their industry. Our research addresses this gap by exploring developers’ nuanced views on the opportunities and challenges of integrating LLMs into professional workflows. By centering the voices of developers, we provide actionable insights into the practical and ethical considerations of adopting LLMs at scale.
5.4.2. Education
Research on the integration of LLMs into education, particularly in computing, has gained attention. Tanay et al. (2024) observed that upper-level computing students using LLMs in software engineering projects improved efficiency in obtaining information and completing tasks, though concerns were raised about potential negative impacts on learning outcomes (Tanay et al., 2024). Essel et al. (2024) found that ChatGPT enhanced undergraduate students’ critical, reflective, and creative thinking skills (Essel et al., 2024). This has sparked discussions on adapting the computer science curriculum to incorporate LLMs. Ozkaya proposed that students should learn to collaborate with LLMs and AI applications, especially for legacy systems (Ozkaya, 2023b). Additionally, Jeuring, Groot, and Keuning highlighted the strong correlation between computational thinking (CT) skills and effective co-development with ChatGPT, suggesting the continued relevance of these skills for both students and future professionals (Jeuring et al., 2023).
While these studies offer insights into how LLMs may enhance student learning, there is limited research on developers’ perspectives regarding the role of LLMs in computing education. Our study addresses this gap by engaging developers in a discussion about how LLMs could reshape education in the field. By focusing on their experiences and views, we provide valuable insights into how LLMs can influence both the learning and teaching of computing skills in professional settings.
6. Threats to validity

### v2

@@ 5.2.1 For Individual Developers: New Competencies and Practices
Key mitigation strategies include: (1) sanitizing generated code by carefully reviewing and cleaning LLM outputs to prevent vulnerabilities; (2) maintaining data integrity by ensuring confidentiality of proprietary code and sensitive data; and (3) upholding security standards by consistently applying established security practices to avoid compromising systems or violating regulations [43].
Cultivating Critical LLM Literacy
Underlying these competencies is critical LLM literacy, discerning when, how, and to what extent to trust LLM outputs. Developers must recognize inherent limitations: hallucinations (6/16), knowledge gaps (5/16), struggles with novel problems (4/16), and inability to provide citations (11/16 desired references). Our findings reveal that LLMs excel at boilerplate generation (10/16), code explanation (11/16), and debugging guidance (10/16), but struggle with requirements gathering (12/16 found no impact), novel problems, and complex reasoning.
Structured courses or workshops can accelerate this learning [77], though experiential practice on non-critical tasks remains valuable for developing the metacognitive awareness that characterizes expert use.
Special Considerations for Solo Developers
Solo developers face unique challenges. Two participants (P9, P10) used LLMs as surrogate team members, or in other words, virtual colleagues for requirements refinement, research, and code review. As P10 mentioned: “I’ve been working solo by myself for a few months now. [ChatGPT] is my junior who [is] trying to help me.” While not substituting for human collaboration, LLMs can enhance independent developers’ self-sufficiency, though vigilance against over-reliance remains essential.
5.2.2 For Organizations: Governance, Guidelines, and Cultural Adaptation
While individual developer competency is necessary, organizational support is equally critical for effective and responsible LLM integration. Our findings empirically confirm that the legal and compliance ambiguities identified in the literature are not theoretical but represent immediate, practical challenges for early adopters. The themes of security and privacy concerns were pervasive, with 13 of 16 participants explicitly expressing concerns over leaking proprietary data. This finding directly operationalizes the literature’s concerns regarding IP attribution and data privacy. Furthermore, our participants were not passive in this regard; they demonstrated active, bottom-up risk mitigation by employing ”security-conscious generalization” in their prompts and actively using tool features to opt out of data training. This developer-led effort was coupled with an explicit organizational void, as 8 of 16 participants highlighted the absence and necessity of guidelines. This result substantiates the ”unresolved compliance issues” cited in prior work, revealing that professionals are currently navigating a high-risk legal landscape largely without the formal governance frameworks they deem necessary.
Establishing Clear Guidelines and Policies
Organizations should establish explicit LLM usage policies addressing four critical dimensions. First, data sharing policies must clarify what information can be shared with external LLM providers. Many participants appropriately generalized prompts to avoid sharing confidential information (13/16), but formalized policies provide clarity and reduce the cognitive burden of case-by-case risk assessment. As P9 emphasized, organizations handling sensitive data require stricter standards: “I think 100% [that companies should have guidelines], especially when you’re dealing with things that are sensitive; for example, financial institution or medical [data].”
Second, code review requirements must specify scrutiny levels for LLM-generated code before production deployment. Fifteen participants modified code before integration, and eight expressed concerns about using unreviewed LLM outputs. Organizations might establish different review tiers: a lightweight review for non-critical components, a rigorous review for security-sensitive code, and a potential prohibition for safety-critical systems.
Third, approved and prohibited use cases should be explicitly defined. Participants noted that LLMs work well for certain tasks (boilerplate generation, debugging guidance) but poorly for others (novel problems, complex architecture, requirements gathering). Organizations can leverage these insights to guide appropriate use, restricting LLM use for security-critical components while encouraging it for routine tasks.
Fourth, documentation and attribution requirements should clarify expectations when LLMs contribute to codebases. Organizations might require developers to note when significant code sections originate from LLMs, supporting future maintenance and enabling tracking of LLM-related issues. Five participants noted that their companies lacked specific guidelines, though existing rules about not sharing sensitive information provided some guardrails. Organizations that establish clear, well-reasoned policies position themselves to capture LLM benefits while mitigating risks.
Providing Technical Infrastructure
Beyond policy, organizations can provide technical infrastructure that supports the effective use of LLMs while maintaining security and control. Participants noted friction when switching between LLM interfaces and development environments. Seamless integration through tools like GitHub Copilot, Amazon Q, and IBM Watsonx Code Assistant reduces workflow disruption and encourages appropriate use by lowering barriers to access.
Organizations with specialized needs might invest in tailored LLM solutions that address concerns about proprietary information while providing contextually-relevant assistance. Retrieval-augmented generation (RAG) [78] enables LLMs to dynamically incorporate project documentation, code repositories, and communication threads into their responses, providing more accurate and context-aware suggestions than generic models. As P13 suggested, tools trained on project documentation could particularly aid developers working with legacy or brownfield codebases, while locally hosted models that never transmit data externally can address the data sovereignty concerns eight participants expressed.
These infrastructure investments serve dual purposes: they support developers in using LLMs effectively while providing organizations with visibility and control over LLM usage patterns. Centralized tools enable monitoring of adoption rates, identification of effective practices worth sharing, and detection of potential misuse or security concerns.
Fostering Cultural and Workflow Adaptation
LLM integration necessitates cultural adaptation beyond policy and infrastructure. Nine participants compared LLMs to team members—pair programmers, junior developers, or assistants—suggesting that teams must reconceptualize workflow and collaboration patterns. As P15 noted, LLMs enable focus on “higher-level stuff that would take more understanding of the infrastructure,” suggesting potential elevation of human work toward more strategic, architectural, and collaborative activities.
This shift affects multiple organizational dimensions. Hiring criteria may evolve, with some participants noting that LLMs might raise the bar for junior positions as basic syntax knowledge becomes less differentiating. Organizations should emphasize problem-solving capabilities, communication skills, and critical thinking rather than focusing narrowly on coding speed. However, nine participants believed the overall job market would remain stable, arguing that software development involves dimensions beyond code writing that LLMs cannot address.
Onboarding practices must evolve to include LLM competency training alongside traditional skills. New developers need guidance on prompt engineering, verification methods, and appropriate use cases, i.e., the competencies our participants developed through months of experimentation. Organizations can accelerate this learning through structured training, mentorship from experienced LLM users, and documented best practices.
Code review norms require recalibration. Reviewers must be vigilant about LLM-generated code, which may appear clean while harboring subtle logical errors, using outdated patterns, or introducing security vulnerabilities. As P11 cautioned: “I think it’s dangerous, right? Like, if out of our work on a team, somebody would just copy-paste ChatGPT [code], you know, I’ll probably be annoyed by it.” Organizations should cultivate a culture where admitting to LLM use is acceptable so that reviewers can apply appropriate scrutiny.
Finally, organizations should foster communities of practice where developers share LLM strategies, lessons learned, and effective approaches. Fourteen participants acknowledged that prompt engineering was essential to their success, yet many developed these skills independently through trial and error. Systematic knowledge-sharing, through documentation, training sessions, or internal forums, can accelerate collective learning and establish organizational standards for quality and security [26].
5.2.3 For Educators: Curriculum and Pedagogy
Computing education faces a fundamental dilemma: LLMs enable students to generate code without understanding it, potentially undermining learning objectives, yet they are ubiquitous tools that students will encounter professionally and are already using regardless of institutional policies. Our participants’ perspectives offer guidance: ten advocated integrating LLMs into education rather than banning them, while six emphasized that fundamental concepts must remain central to curricula. This balance, i.e., embracing LLMs as learning tools while ensuring deep conceptual understanding, defines the educational challenge.
Our findings directly address the precise literature gap regarding practitioners’ perceived effects of LLMs on education. While prior work has focused on the student-centric risks of over-reliance, our study captures the missing practitioner perspective, confirming that early adopters see an urgent need for reshaping computing education (a theme endorsed by 9 of 16 participants). Our participants’ views operationalize the abstract concerns from the literature: they advocate for curricula to de-emphasize the ”commodity skills” (e.g., syntax, boilerplate) that LLMs now automate, and instead intensify focus on ”fundamental concepts” like problem-solving, algorithms, and critical thinking. Furthermore, our results extend this discussion beyond just what to de-emphasize by identifying the specific new skills practitioners demand.
The Impracticality and Inequity of Prohibition
Three participants explicitly noted the impracticality of prohibiting LLM use. As P2 stated: “You cannot run away from it because everyone has access […] Students are going to use it, no matter [what].” Prohibition creates multiple problems. First, it is unenforceable: students can easily access LLMs on personal devices, making bans ineffective without intrusive monitoring. Second, it disadvantages compliant students: those who follow bans are left unprepared for professional practice where LLM use is standard, while peers who violate bans gain unfair advantages and valuable experience. Third, it misses educational opportunities: LLMs can democratize access to educational support, benefiting students who lack resources for tutoring, as P6 noted: “[ChatGPT] has lowered the barrier to entry in terms of coding.”
Rather than fighting inevitable technology adoption, educators should channel it productively. The challenge is how to design curricula that leverage LLM benefits while ensuring genuine learning, acknowledging that some traditional exercises lose pedagogical value when LLMs can complete them instantly.
Curriculum Content: Fundamentals and LLM Literacy
Effective curricula must balance two objectives: preserving fundamental knowledge and cultivating LLM literacy. Six participants emphasized that foundational concepts, such as data structures, algorithms, software architecture, and problem-solving, remain critical. As P1 stressed, universities “should be doing more of what they should have been doing in the first place, which is not necessarily focusing on specific implementations […] but rather on the bigger, more difficult-to-grasp ideas that have to do with architecting software.”
These fundamentals are not threatened by LLMs; they are essential for using LLMs effectively. Our findings reveal that developers need a deep understanding to evaluate LLM outputs critically, identify subtle errors, make architectural decisions, and formulate problems effectively. Students who lack foundational knowledge cannot assess whether LLM-generated code is correct, efficient, or appropriate. As P5 emphasized, LLMs require “foundational knowledge of your data structures, algorithms, how to design code databases, SDLC, all of those different core topics.” The strongest argument for teaching fundamentals is that LLMs make them more important—students need robust conceptual understanding to serve as critical evaluators of generated code.
Simultaneously, curricula should explicitly teach LLM literacy, encompassing an understanding of how LLMs work, their capabilities and limitations across various tasks, prompt engineering techniques, and critical evaluation practices. This is analogous to teaching “how to Google effectively,” as P13 noted—a meta-skill about effectively using available tools. Four participants emphasized the importance of teaching prompt engineering, recognizing it as a valuable professional competency.
LLM literacy curricula might include modules on: (1) how LLMs work—basic understanding of training processes and probabilistic nature to manage expectations; (2) capabilities across the SDLC—understanding where LLMs excel (implementation, debugging) and struggle (requirements, novel problems); (3) prompt engineering practice—developing the six strategies our participants employed; (4) verification techniques—teaching the five methods our participants used; and (5) security and ethical considerations—understanding data privacy, code provenance, and responsible use. Importantly, LLM literacy should be integrated throughout the curriculum, rather than taught as a separate course, to demonstrate its appropriate use in various contexts.
Pedagogical Approaches: Assignment Redesign and Authentic Assessment
Pedagogy must evolve alongside curriculum. Two participants suggested redesigning assignments to require genuine understanding even when LLMs are available. As P2 stated: “You can design the assignments in such a way that even if [students] use [ChatGPT], they need to understand the concepts.”
Several pedagogical strategies can achieve this goal. First, emphasize explanation and justification over code submission. Assignments might ask students to implement solutions, then explain design decisions, analyze trade-offs, and justify choices. LLMs can generate implementations, but explaining why requires understanding. Second, use process-oriented assessment. Require students to document their development process, including initial problem understanding, challenges encountered, how they utilized LLMs, and what they learned. This shifts evaluation from product to process, making LLM use transparent and assessable.
Third, employ oral examinations or live coding to assess understanding. If students cannot explain code when questioned or cannot solve similar problems without assistance, this reveals insufficient learning. Fourth, design integration-focused assignments that require students to combine LLM outputs with original work, demonstrating synthesis skills. Fifth, create assignments requiring novel problem-solving where LLMs struggle. Our findings reveal that LLMs struggle with novel ideas and complex logical reasoning (4/16 noted struggles), so assignments that require creative problem-solving emphasize higher-order thinking.
Finally, teach students to use LLMs as learning tools rather than solution generators. Educators can model appropriate use: asking LLMs to explain concepts, generate practice problems, or identify errors in student code. When students view LLMs as tutors rather than assignment completers, they engage more deeply with the material. Ten participants advocated encouraging LLM use as learning tools, recognizing their potential to support rather than subvert education when used intentionally.
Addressing the Learning Impediment Concern
Four participants raised concerns about LLMs potentially impeding learning, particularly for those who may lack expertise to evaluate generated code. This risk is real: students might use LLMs to complete assignments without developing genuine understanding. However, this risk exists with any educational technology—the solution is not prohibition but deliberate pedagogical design. This requires: (1) scaffolded introduction where students first learn fundamentals, then gradually incorporate LLM use; (2) explicit instruction on when and how to use LLMs effectively for learning; (3) assessment redesign that rewards understanding over code production; and (4) metacognitive development where students recognize when LLM use is helping versus hindering their growth.
As P5 articulated: “If universities are trying to prepare students for work in the industry, I would say that exposing the students to how to properly use ChatGPT to create impactful works in code and program would certainly be beneficial in the professional corporate setting. But, you still need to have the foundational knowledge.” This balanced perspective—embracing LLMs as professional tools while maintaining rigorous foundational education—should guide curricular reform.
5.3 The Evolving LLM Landscape: Progress and Persistent Challenges
Since our interviews were conducted during the Spring and Summer of 2023, the landscape of LLMs has notably evolved. Understanding these changes is crucial for evaluating the continued relevance of our findings and identifying which limitations have been addressed versus those that persist.
5.3.1 Persistent Limitations and Implications for Longevity
Despite impressive advances in commercially available large language models from major AI companies such as OpenAI, Google, and Anthropic, which include massively expanded context window sizes, reasoning capabilities, knowledge currency, and performance, several categories of limitations still persist, as empirically validated by recent 2025 studies. Technical limitations include: (1) struggles with truly novel or highly specialized problems where training data is sparse; (2) maintaining perfect consistency across very long, complex interactions; (3) providing citations or transparency about information sources—eleven participants desired references, and this persists as LLMs rely primarily on parametric knowledge; and (4) avoiding subtle semantic errors that appear syntactically correct, requiring continued human vigilance [26, 79, 25, 73].
Moreover, while flagship models have addressed many limitations, smaller or locally deployable models that many organizations favor for privacy or cost-efficiency continue to exhibit more pronounced constraints. Organizations that cannot utilize cloud-based LLMs due to data sovereignty concerns face distinct trade-offs, making our findings on verification practices and security-conscious integration particularly relevant.
For organizations prioritizing the scale and efficiency that large language models can provide, the literature reveals a major constraint in LLMs’ ability to handle large, repository-level contexts, systematically quantified by a new generation of benchmarks [79, 73]. The findings from ML-BENCH are particularly revealing: they demonstrate that even reasoning models have a low absolute success rate on these repository-scale tasks, which are the norm in modern corporate contexts. In direct response to this ”context-gap,” a new paradigm of ”Agentic Software Engineering” is emerging, which moves beyond simple code generation toward achieving complex, goal-oriented software engineering objectives [80].
More fundamentally, many limitations we documented are sociotechnical rather than purely technical, rendering them resistant to model improvements alone. LLMs’ potential to impede learning (4/16 expressed concern) stems from how they are used rather than their technical capabilities—even perfect code generation could undermine learning if students do not engage in problem-solving. Similarly, LLMs’ inability to replace human collaboration, stakeholder communication, and domain expertise (12/16 noted this for requirements gathering) reflects the inherently human nature of these activities. Security and privacy concerns around data sharing (8/16) remain salient as long as LLMs are trained on user inputs or hosted externally. The expertise gap between novice and expert users persists—our participants developed sophisticated strategies through intensive experimentation, while later adopters still face substantial learning curves.
The relevance of our findings extends beyond the specific models used by our participants for two reasons. First, the patterns we documented reflect fundamental characteristics of software work and LLM-mediated development rather than transient model weaknesses. The phase-dependent adoption pattern—high utility in implementation and debugging, with minimal impact on requirements gathering—persists because it reflects the nature of these activities (technical versus collaborative, well-defined versus ambiguous) rather than limitations of the model. The productivity-quality paradox, i.e., time savings coexisting with high code discard rates, reveals how LLMs shift cognitive work rather than eliminate it, a pattern that remains relevant as long as human oversight is necessary.
The competencies we identified—prompt engineering, critical evaluation, security-conscious integration—remain essential across model generations. While specific strategies may evolve, the underlying skill of formulating problems effectively for LLM consumption persists. The verification methods our participants employed remain necessary as long as LLMs can generate plausible but incorrect code.
Second, early adopters establish practices that become institutionalized as technologies mature. The six prompt engineering strategies that our participants developed through experimentation are now featured in formal training materials. Their verification workflows inform organizational policies. Their security concerns shape regulatory discussions and vendor features. By documenting these emergent practices during the formation stage, our study captures not only outcomes but also the reasoning and learning processes that shaped effective approaches—insights that remain valuable for understanding how to utilize LLMs effectively, even as the tools evolve.
Future developers integrating new LLM capabilities or organizations adopting LLM technologies will face similar challenges: learning what works through experimentation, developing verification strategies, balancing productivity with quality, and establishing appropriate governance. Our participants’ experiences provide a valuable template for navigating these challenges, making our findings a roadmap for LLM integration rather than merely a snapshot of 2023 tool capabilities. As P6 anticipated, “The fundamentals remain the same.” While LLM capabilities improve, the human work of problem formulation, critical evaluation, collaborative communication, and responsible integration persists.
6 Threats to validity